# Exploratory Randomization for Discrete-Time Linear Exponential Quadratic Gaussian (LEQG) Problem

### v1

@@ 3.2 Energy-Entropy Duality for the Randomized LEQG Problem
This section aims to establish the relation between the risk-sensitive problem and the dual risk-neutral penalized problem. For this purpose, we introduce on the measurable space $\left(\Omega,\mathcal{F},\left(\mathcal{F}\right)_{t=0,\ldots,T}\right)$ a probability measure $\mathbb{P}^{\bar{\gamma},\bar{\eta}}$, different from $\mathbb{P}$, and parametrized by two vectors $\bar{\gamma}\in\mathbb{R}^{d_{x}},\bar{\eta}\in\mathbb{R}^{d_{u}}$.
Under measure $\mathbb{P}^{\bar{\gamma},\bar{\eta}}$, the state noise $\bar{w}_{t}$ is i.i.d. Gaussian with mean $\bar{\gamma}_{t}$, which may depend on time, and covariance $\Lambda_{t}$, namely $w^{\bar{\gamma}}_{t}\sim\mathcal{N}\left(\bar{\gamma}_{t},\Lambda_{t}\right)$. Furthermore, the noise term $\bar{\eta}$ affects the randomized control $u_{t}=\bar{u}_{t}+v^{\bar{\eta}}_{t}$ via its noise $v^{\bar{\eta}}_{t}\sim\mathcal{N}\left(\bar{\eta}_{t},\Xi_{t}\right)$, that is, the control noises are i.i.d. Gaussian with mean $\bar{\eta}_{t}$ and covariance $\Xi_{t}$. We represent the randomized control as the measure $\pi\left(du;\bar{u}_{t},\bar{\eta}_{t}\right)\sim\mathcal{N}\left(\bar{u}_{t}+\bar{\eta}_{t},\Xi_{t}\right)$.
@@
$\frac{d\mathbb{P}^{\bar{\gamma},\bar{\eta}}}{d\lambda}=$ $\prod_{t=0}^{T-1}\left(\frac{1}{\sqrt{(2\pi)^{d_{x}}|\Lambda_{t}|}}\exp\left\{-\frac{1}{2}\left(w_{t}-\bar{\gamma}_{t}\right)^{\prime}\Lambda^{-1}_{t}\left(w_{t}-\bar{\gamma}_{t}\right)\right\}\right)$
×$\times\left(\frac{1}{\sqrt{(2\pi)^{d_{u}}|\Xi_{t}|}}\exp\left\{-\frac{1}{2}\left(v_{t}-\bar{\eta}_{t}\right)^{\prime}\Xi_{t}^{-1}\left(v_{t}-\bar{\eta}_{t}\right)\right\}\right).$ (3.5)
The expression for $\frac{d\mathbb{P}}{d\lambda}$ takes the same form, but with $\bar{\gamma}=0$ and $\bar{\eta}=0$. Therefore,
[…]
$=$ $\prod_{t=0}^{T-1}\exp\left\{\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}w_{t}-\frac{1}{2}\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}\right\}\exp\left\{\bar{\eta}^{\prime}\Xi_{t}^{-1}v_{t}-\frac{1}{2}\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right\}$ (3.6)
@@
$D_{\text{KL}}\left(\mathbb{P}^{\bar{\gamma},\bar{\eta}}\|\mathbb{P}\right)$
@@
$=$ $\sum_{t=0}^{T-1}\mathbf{E}^{\bar{\gamma},\bar{\eta}}\left[\left\{\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}w_{t}-\frac{1}{2}\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}\right\}+\left\{\bar{\eta}^{\prime}\Xi_{t}^{-1}v_{t}-\frac{1}{2}\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right\}\right]$
$=$ $\sum_{t=0}^{T-1}\left\{\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\mathbf{E}^{\bar{\gamma},\bar{\eta}}\left[w_{t}\right]-\frac{1}{2}\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}\right\}+\left\{\bar{\eta}^{\prime}\Xi_{t}^{-1}\mathbf{E}^{\bar{\gamma},\bar{\eta}}\left[v_{t}\right]-\frac{1}{2}\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right\}$
$=$ $\sum_{t=0}^{T-1}\left\{\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}-\frac{1}{2}\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}\right\}+\left\{\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}-\frac{1}{2}\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right\}$
$=$ $\frac{1}{2}\sum_{t=0}^{T-1}\left(\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}+\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right)$ (3.7)
@@
for $t=0,\ldots,T-1$. Furthermore, the reward function at (2.2) can be expressed as:
$G_{T}:=$ $\sum_{t=0}^{T-1}\left[x_{t}^{\prime}Mx_{t}+\int u_{t}^{\prime}N_{t}u_{t}\pi\left(du;\bar{u}_{t},\bar{\eta}_{t}\right)+\int u_{t}^{\prime}Qx_{t}\pi\left(du;\bar{u}_{t},\bar{\eta}_{t}\right)+x_{t}^{\prime}m+\int u_{t}^{\prime}n\pi\left(du;\bar{u}_{t},\bar{\eta}_{t}\right)\right]$
$+x_{T}^{\prime}M_{T}x_{T}+x_{T}^{\prime}m_{T}$
$=$ $\sum_{t=0}^{T-1}\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi N_{t}\right)+\left(\bar{u}_{t}+\bar{\eta}\right)^{\prime}N_{t}\left(\bar{u}_{t}+\bar{\eta}\right)+\left(\bar{u}_{t}+\bar{\eta}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(\bar{u}_{t}+\bar{\eta}\right)^{\prime}n\right]$
$+x_{T}^{\prime}M_{T}x_{T}+x_{T}^{\prime}m_{T}.$ (3.9)
@@
$\ln I(\bar{u};T,\theta)=\sup_{\bar{\gamma},\bar{\eta}}\mathbf{E}^{\bar{\gamma},\bar{\eta}}\left[\theta G_{T}-\frac{1}{2}\sum_{t=0}^{T-1}\left(\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}+\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right)\right].$ (3.10)
Taking the infimum over $\bar{u}$, we have
@@
where the first equivalence follows from Lemma 5.3.1 in Meneghini, (1994).
@@
We can thus interpret the energy-entropy duality’s $\inf\sup$ as a two-player game against Nature. The agent applies control $u$ to minimize the expectation while Nature (via the duality) applies control $\bar{\nu}:=(\bar{\gamma}^{\prime}\bar{\eta}^{\prime})^{\prime}$ to maximize it. Notice that we subtract the entropy from the cost criterion $I(u;T,\theta)$ to be minimized in $\bar{u}$; on the other hand we maximize over the entropy parameters $\bar{\gamma}$ and $\bar{\eta}$. In Wang et al., (2020), the authors simply add the randomized control entropy to the criterion to be minimized with respect to the originally given control as a reasonable way to proceed but without further explanation. Their entropy is not parametrized and therefore their problem does not become a stochastic differential game. By contrast, our setup results from the energy-entropy duality, justofying the presence of an additive entropy penalization.
@@ 3.3 Solving the Risk-Neutral Penalized Game Problem
To solve the stochastic game with optimal value $V(T;\theta)$ as in (3.12) we shall apply the Dynamic Programming Principle (DPP). To this effect let $V_{t}(x_{t})$ be the optimal value of the game at the generic time $t\quad(t=0,\cdots,T)$ when the controlled state takes the value $x_{t}$. The risk sensitivity parameter $\theta$ is supposed to be fixed. Using (3.12) that expresses the global optimal value of the game as well as (3.2) that expresses the global cost $G_{T}$, the DPP lets us express $V_{t}$ recursively as
$V_{T}(x_{T})=$ $\theta\,\left[x_{T}^{\prime}M_{T}x_{T}+x_{T}^{\prime}m_{T}\right]$
@@
$V_{t}(x_{t})=$ $\inf_{\bar{u}}\sup_{\bar{\gamma},\bar{\eta}}\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left\{\theta\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi N_{t}\right)+\left(\bar{u}+\bar{\eta}\right)^{\prime}N_{t}\left(\bar{u}+\bar{\eta}\right)+\left(\bar{u}+\bar{\eta}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m\right.\right.$
$\left.\left.+\left(\bar{u}+\bar{\eta}\right)^{\prime}n\right]-\frac{1}{2}\left(\bar{\gamma}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}+\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}\right)+V_{t+1}(x_{t+1})\right\}$
$=$ $\inf_{\bar{u}}\sup_{\bar{\gamma},\bar{\eta}}\left\{\theta\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi N_{t}\right)+\left(\bar{u}+\bar{\eta}\right)^{\prime}N_{t}\left(\bar{u}+\bar{\eta}\right)+\left(\bar{u}+\bar{\eta}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m\right.\right.$
$\left.\left.+\left(\bar{u}+\bar{\eta}\right)^{\prime}n\right]-\frac{1}{2}\left(\bar{\gamma}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}+\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}\right)+\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[V_{t+1}(x_{t+1})\right]\right\},$ (3.15)
@@
$\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[V_{t+1}(x_{t+1})\right]$
$=$ $\frac{1}{2}\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}\left(a+Ax_{t}+B\bar{u}_{t}\right)+\frac{1}{2}\bar{\gamma}_{t}^{\prime}P_{t+1}\bar{\gamma}_{t}+\frac{1}{2}\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}B\bar{\eta}_{t}+\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}\bar{\gamma}_{t}$
$+$ $\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}p_{t+1}+\bar{\gamma}_{t}^{\prime}p_{t+1}+\bar{\eta}_{t}^{\prime}B^{\prime}p_{t+1}+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}\bar{\gamma}_{t}+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}B\bar{\eta}_{t}$
$+$ $\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}.$ (3.17)
@@
$\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[V_{t+1}(x_{t+1})\right]$
$=$ $\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[\frac{1}{2}\left(a+Ax_{t}+B\left(\bar{u}_{t}+v^{\bar{\eta}}_{t}\right)+w^{\bar{\gamma}}_{t}\right)^{\prime}P_{t+1}\left(a+Ax_{t}+B\left(\bar{u}_{t}+v^{\bar{\eta}}_{t}\right)+w^{\bar{\gamma}}_{t}\right)\right.$
@@
$=$ $\frac{1}{2}\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}\left(a+Ax_{t}+B\bar{u}_{t}\right)\right]+\frac{1}{2}\underbrace{\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[(v^{\bar{\eta}}_{t})^{\prime}B^{\prime}P_{t+1}Bv^{\bar{\eta}}_{t}\right]}_{=\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}B\bar{\eta}_{t}}+\frac{1}{2}\underbrace{\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[(w^{\bar{\gamma}}_{t})^{\prime}P_{t+1}w^{\bar{\gamma}}_{t}\right]}_{=\text{tr}\left(\Lambda_{t}P_{t+1}\right)+\bar{\gamma}_{t}^{\prime}P_{t+1}\bar{\gamma}_{t}}$
$+\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}Bv^{\bar{\eta}}_{t}\right]+\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}w^{\bar{\gamma}}_{t}\right]+\underbrace{\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[(v^{\bar{\eta}}_{t})^{\prime}B^{\prime}P_{t+1}w^{\bar{\gamma}}_{t}\right]}_{=\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}\bar{\gamma}_{t}}$

### v2

@@ 3.2 Energy-Entropy Duality for the Randomized LEQG Problem
This section aims to establish the relation between the risk-sensitive problem and the dual risk-neutral penalized problem. For this purpose, we introduce on the measurable space $\left(\Omega,\mathcal{F},\left(\mathcal{F}\right)_{t=0,\ldots,T}\right)$ a probability measure $\mathbb{P}^{\bar{\gamma},\bar{\eta}}$, different from $\mathbb{P}$, and parametrized by two processes $\bar{\gamma}=(\bar{\gamma}_{t})_{t=0,\cdots,T}\in\mathbb{R}^{d_{x}},\bar{\eta}=(\bar{\eta}_{t})_{t=0,\cdots,T}\in\mathbb{R}^{d_{u}}$.
Under measure $\mathbb{P}^{\bar{\gamma},\bar{\eta}}$, the state noise $w^{\bar{\gamma}}_{t}$ is i.i.d. Gaussian with mean $\bar{\gamma}_{t}$, which may depend on time, and covariance $\Lambda_{t}$, namely $w^{\bar{\gamma}}_{t}\sim\mathcal{N}\left(\bar{\gamma}_{t},\Lambda_{t}\right)$. Furthermore, the process $\bar{\eta}$ affects the randomized control $u_{t}=\bar{u}_{t}+v^{\bar{\eta}}_{t}$ via its noise $v^{\bar{\eta}}_{t}\sim\mathcal{N}\left(\bar{\eta}_{t},\Xi_{t}\right)$, that is, the control noises are i.i.d. Gaussian with mean $\bar{\eta}_{t}$ and covariance $\Xi_{t}$. We represent the randomized control as the measure $\pi\left(du;\bar{u}_{t},\bar{\eta}_{t}\right)\sim\mathcal{N}\left(\bar{u}_{t}+\bar{\eta}_{t},\Xi_{t}\right)$.
@@
$\frac{d\mathbb{P}^{\bar{\gamma},\bar{\eta}}}{d\lambda}=$ $\prod_{t=0}^{T-1}\left(\frac{1}{\sqrt{(2\pi)^{d_{x}}|\Lambda_{t}|}}\exp\left\{-\frac{1}{2}\left(w_{t}-\bar{\gamma}_{t}\right)^{\prime}\Lambda^{-1}_{t}\left(w_{t}-\bar{\gamma}_{t}\right)\right\}\right)$ (3.5)
$\times\left(\frac{1}{\sqrt{(2\pi)^{d_{u}}|\Xi_{t}|}}\exp\left\{-\frac{1}{2}\left(v_{t}-\bar{\eta}_{t}\right)^{\prime}\Xi_{t}^{-1}\left(v_{t}-\bar{\eta}_{t}\right)\right\}\right).$
The expression for $\frac{d\mathbb{P}}{d\lambda}$ takes the same form, but with $\bar{\gamma}_{t}=0$ and $\bar{\eta}_{t}=0$. Therefore,
$\frac{d\mathbb{P}^{\bar{\gamma},\bar{\eta}}}{d\mathbb{P}}=$ $\prod_{t=0}^{T-1}\frac{\left(\cancel{\frac{1}{\sqrt{(2\pi)^{d_{x}}|\Lambda_{t}|}}}\exp\left\{-\frac{1}{2}\left(w_{t}-\bar{\gamma}_{t}\right)^{\prime}\Lambda^{-1}_{t}\left(w_{t}-\bar{\gamma}_{t}\right)\right\}\right)}{\left(\cancel{\frac{1}{\sqrt{(2\pi)^{d_{x}}|\Lambda_{t}|}}}\exp\left\{-\frac{1}{2}w_{t}^{\prime}\Lambda^{-1}_{t}w_{t}\right\}\right)}$ (3.6)
$\times\frac{\left(\cancel{\frac{1}{\sqrt{(2\pi)^{d_{u}}|\Xi_{t}|}}}\exp\left\{-\frac{1}{2}\left(v_{t}-\bar{\eta}_{t}\right)^{\prime}\Xi_{t}^{-1}\left(v_{t}-\bar{\eta}_{t}\right)\right\}\right)}{\left(\cancel{\frac{1}{\sqrt{(2\pi)^{d_{u}}|\Xi_{t}|}}}\exp\left\{-\frac{1}{2}v_{t}^{\prime}\Xi_{t}^{-1}v_{t}\right\}\right)}$
$=$ $\prod_{t=0}^{T-1}\exp\left\{\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}w_{t}-\frac{1}{2}\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}\right\}\exp\left\{\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}v_{t}-\frac{1}{2}\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right\}$
@@
$D_{\text{KL}}\left(\mathbb{P}^{\bar{\gamma},\bar{\eta}}\|\mathbb{P}\right)$ (3.7)
@@
$=$ $\sum_{t=0}^{T-1}\mathbf{E}^{\bar{\gamma},\bar{\eta}}\left[\left\{\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}w_{t}-\frac{1}{2}\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}\right\}+\left\{\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}v_{t}-\frac{1}{2}\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right\}\right]$
$=$ $\sum_{t=0}^{T-1}\left\{\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\mathbf{E}^{\bar{\gamma},\bar{\eta}}\left[w_{t}\right]-\frac{1}{2}\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}\right\}+\left\{\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\mathbf{E}^{\bar{\gamma},\bar{\eta}}\left[v_{t}\right]-\frac{1}{2}\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right\}$
$=$ $\sum_{t=0}^{T-1}\left\{\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}-\frac{1}{2}\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}\right\}+\left\{\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}-\frac{1}{2}\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right\}$
$=$ $\frac{1}{2}\sum_{t=0}^{T-1}\left(\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}+\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right)$
@@
for $t=0,\ldots,T-1$. Furthermore, the cost function at (2.6) can be expressed as:
$G_{T}:=$ $\sum_{t=0}^{T-1}\left[x_{t}^{\prime}Mx_{t}+\int u_{t}^{\prime}N_{t}u_{t}\pi\left(du;\bar{u}_{t},\bar{\eta}_{t}\right)+\int u_{t}^{\prime}Qx_{t}\pi\left(du;\bar{u}_{t},\bar{\eta}_{t}\right)\right$. (3.9)
$\left.+x_{t}^{\prime}m+\int u_{t}^{\prime}n\pi\left(du;\bar{u}_{t},\bar{\eta}_{t}\right)\right]+x_{T}^{\prime}M_{T}x_{T}+x_{T}^{\prime}m_{T}$
$=$ $\sum_{t=0}^{T-1}\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi_{t}N_{t}\right)+\left(\bar{u}_{t}+\bar{\eta}_{t}\right)^{\prime}N_{t}\left(\bar{u}_{t}+\bar{\eta}_{t}\right)\right$.
$\left.+\left(\bar{u}_{t}+\bar{\eta}_{t}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(\bar{u}_{t}+\bar{\eta}_{t}\right)^{\prime}n\right]+x_{T}^{\prime}M_{T}x_{T}+x_{T}^{\prime}m_{T}.$
@@
$\ln I(\bar{u};T,\theta)=\sup_{\bar{\gamma},\bar{\eta}}\mathbf{E}^{\bar{\gamma},\bar{\eta}}\left[\theta G_{T}-\frac{1}{2}\sum_{t=0}^{T-1}\left(\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}+\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right)\right],$ (3.10)
where $I$ is the criterion defined at (2.4).
Taking the infimum over $\bar{u}$ to minimize the criterion $I$, we have
@@
where the first equivalence follows from Lemma 5.3.1 in Meneghini (1994).
@@
We can thus interpret the energy-entropy duality’s $\inf\sup$ as a two-player game against Nature. The agent applies control $u$ to minimize the expectation while Nature (via the duality) applies control $\bar{\nu}:=\begin{pmatrix}\bar{\gamma}^{\prime}&\bar{\eta}^{\prime}\end{pmatrix}^{\prime}$ to maximize it. Notice that we subtract the entropy from the criterion $I(u;T,\theta)$ to be minimized in $\bar{u}$; on the other hand, we maximize over the entropy parameters $\bar{\gamma}$ and $\bar{\eta}$. In Wang et al. (2020), the authors simply add the randomized control entropy to the criterion to be minimized with respect to the originally given control as a reasonable way to proceed but without further explanation. Their entropy is not parametrized and therefore their problem does not become a stochastic differential game. By contrast, our setup results from the energy-entropy duality, justifying the presence of an additive entropy penalization.
@@ 3.3 Solving the Risk-Neutral Penalized Game Problem
To solve the stochastic game with optimal value $V(T;\theta)$ as in (3.12) we shall apply the Dynamic Programming Principle (DPP). To this effect, let $V_{t}(x_{t})$ be the optimal value of the game at the generic time $t\quad(t=0,\cdots,T)$ when the controlled state takes the value $x_{t}$. The risk sensitivity parameter $\theta$ is supposed to be fixed. Using (3.12) that expresses the global optimal value of the game as well as (3.9) that expresses the cost $G_{T}$, the DPP lets us express $V_{t}$ recursively as
$V_{T}(x_{T})=$ $\theta\,\left[x_{T}^{\prime}M_{T}x_{T}+x_{T}^{\prime}m_{T}\right]$ (3.14)
and, for $t=T-1,\cdots,1,0$,
@@
$\left.\left.+\left(\bar{u}+\bar{\eta}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(\bar{u}+\bar{\eta}\right)^{\prime}n\right]\right$.
$\left.-\frac{1}{2}\left(\bar{\gamma}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}+\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}\right)+V_{t+1}(x_{t+1})\right\}$
[…]
$\left.\left.+\left(\bar{u}+\bar{\eta}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(\bar{u}+\bar{\eta}\right)^{\prime}n\right]\right$.
$\left.-\frac{1}{2}\left(\bar{\gamma}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}+\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}\right)+\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[V_{t+1}(x_{t+1})\right]\right\},$
@@
$\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[V_{t+1}(x_{t+1})\right]$ (3.17)
$=$ $\frac{1}{2}\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}\left(a+Ax_{t}+B\bar{u}_{t}\right)+\frac{1}{2}\bar{\gamma}_{t}^{\prime}P_{t+1}\bar{\gamma}_{t}$
$+\frac{1}{2}\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}B\bar{\eta}_{t}+\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}\bar{\gamma}_{t}+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}p_{t+1}+\bar{\gamma}_{t}^{\prime}p_{t+1}$
$+\bar{\eta}_{t}^{\prime}B^{\prime}p_{t+1}+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}\bar{\gamma}_{t}+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}B\bar{\eta}_{t}$
$+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}.$
@@
$\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[V_{t+1}(x_{t+1})\right]=$ $\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[\frac{1}{2}\left(a+Ax_{t}+B\left(\bar{u}_{t}+v^{\bar{\eta}}_{t}\right)+w^{\bar{\gamma}}_{t}\right)^{\prime}P_{t+1}\left(a+Ax_{t}+B\left(\bar{u}_{t}+v^{\bar{\eta}}_{t}\right)+w^{\bar{\gamma}}_{t}\right)\right$.
@@
$=$ $\frac{1}{2}\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}\left(a+Ax_{t}+B\bar{u}_{t}\right)\right]+\frac{1}{2}\underbrace{\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[(v^{\bar{\eta}}_{t})^{\prime}B^{\prime}P_{t+1}Bv^{\bar{\eta}}_{t}\right]}_{=\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}B\bar{\eta}_{t}}$
$+\frac{1}{2}\underbrace{\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[(w^{\bar{\gamma}}_{t})^{\prime}P_{t+1}w^{\bar{\gamma}}_{t}\right]}_{=\text{tr}\left(\Lambda_{t}P_{t+1}\right)+\bar{\gamma}_{t}^{\prime}P_{t+1}\bar{\gamma}_{t}}+\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}Bv^{\bar{\eta}}_{t}\right]$
$+\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}w^{\bar{\gamma}}_{t}\right]+\underbrace{\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[(v^{\bar{\eta}}_{t})^{\prime}B^{\prime}P_{t+1}w^{\bar{\gamma}}_{t}\right]}_{=\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}\bar{\gamma}_{t}}$